OpenStudy (candy13106):

24 miles over hours = blank feet over secondes

OpenStudy (candy13106):

plz help anyone plz

OpenStudy (anonymous):

The conversion from miles per hour to feet per second is: \[\frac{x\phantom{.}feet}{second}=\frac{5280}{3600}\times \frac{y\phantom{.}miles}{hour}\]

OpenStudy (candy13106):

so thats the answer right

OpenStudy (candy13106):

u there

OpenStudy (candy13106):

hello plz i got to go soon man come on

OpenStudy (candy13106):

plz hurry

OpenStudy (anonymous):

Well like if you had. 24 miles\hour, then \[24\times\frac{5280}{3600}=x\] Where x is feet per second

OpenStudy (candy13106):

so thats the answer right

OpenStudy (candy13106):

and wat does x = to

OpenStudy (candy13106):

huryr

OpenStudy (candy13106):

goshyyyyyyyyyyy

OpenStudy (candy13106):

come on plz zzzzzzzzzzzzz

OpenStudy (anonymous):

yeah. x=35.2

OpenStudy (candy13106):

ok thx a lot bye

OpenStudy (candy13106):

c u tomorow if u going to be on